# Measurement-device agnostic quantum tomography

### v4

@@ V Special case: multiplicative error
$\theta_{j}^{\prime}=(1+\delta)\theta_{j},$ (6)
$\phi_{j}^{\prime}=(1+\epsilon)\phi_{j}.$ (7)
@@
Because parameters $(\delta,\epsilon)$ are shared among all projection operators, the optimum search simplifies from 12 parameters to just two. In Section S2 of the Supplementary material, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable. The advantage is better performance when compared to global optimization and also the smaller probe state set. Another advantage is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements.
@@ VI Experiment: polarization state tomography
We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario, an example of a situation with multiplicative errors. At the end of the process, we will improve our knowledge of the waveplate retardance and consequently mitigate the reconstruction artifact.
The experimental setup is depicted in Fig. 1(a). We used a continuous-wave 810 nm thermally-stabilized fiber-coupled laser diode (LD). The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism (GT). The waveplates HWP1 and QWP1 effectively control the prepared polarization state. The following operator describes the action of a waveplate:
@@
where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise. This relation holds for wave plates that have perfect retardance. The deviation from the ideal retardance will introduce errors in the preparation. We assumed the perfect retardance in the first step of the experiment as our method does not require exact knowledge of probe states.
@@
To demonstrate the tolerance of the method to a unitary operation between preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction. Because this unitary operation can be in principle unknown, we restrict ourselves from using any information about QWPX in the calibration.
@@
The states were analyzed using half-waveplate and quarter-wave (HWP2, QWP2) and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer (BD). The waveplate settings for the Pauli tomography are listed in Tab. S2 of the Supplementary Material. All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of ±$\pm 30$ mdeg.
@@
We used reverse-biased PIN photodiodes (PD) to measure optical intensity at both outputs of the beam-displacer. We normalized the readings using the sum of two signals to eliminate laser power fluctuation.
@@
The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually slightly deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta$ and $\epsilon$, respectively. These deviations are initially unknown to us, and we will use the presented method to estimate their values.
@@
We use eight probe states and perform the Pauli tomography for each. Initially, we assume zero deviations, $\delta=\epsilon=0$, and observe the purity modulation $\Delta P=0.0572(3)$. The number in parenthesis is one standard deviation at the last significant digit, e.g. $0.0572(3)=(572\pm 3)\cdot 10^{-4}$. The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision. The shot noise is negligible at the power level of our laser, compared to the technical noise. Then, we vary parameters $\delta$ and $\epsilon$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$. We found the optimal parameters $\delta=4.5(2)$ deg, $\epsilon=-1.30(4)$ deg.
@@
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in state preparation. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation from 0.0700(3) to 0.0144(2) to find the optimal parameters $\tilde{\delta}=4.53(6)$ deg and $\tilde{\epsilon}=-3.58(3)$ deg.
@@
The calibration step revealed the retardance deviation of the wave plates used for preparation and projection. To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators. However, we take one extra step which allows us to prepare the states with greater precision and control the measurement operators precisely. We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation. Since Eq. 10 does not hold anymore, we used numerical optimization to find the desired wave plate settings ($x,y$):
@@
As a reference, we first assumed perfect waveplates and performed state tomography for each test state. We indeed observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$. This modulation indicates reconstruction artifacts. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.96688(4)$. These fidelity values indicate imprecise preparation.
In contrast, when we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction and preparation, the minimum purity increased to $0.9830(3)$ and the minimum fidelity to $F_{\min}=0.9907(2)$. The information gained in the calibration clearly improved the reconstruction.
@@ VII Path-encoded qubits on a photonic chip
Figure 3: Application of the method for on-chip photonic path qubit. (a) Projection apparatus for a path qubit. Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers. These shifts determine the measurement projector. The photons in the upper output path are detected with a single-photon detector (SPAD). (b) 1D-section of 6D function $\Delta P(c_{1},\dots,c_{6})$. The value of $\Delta P$ at the initial assumption of $c_{j}$ is $1.8\cdot 10^{-2}$ on the ensemble of probe states. Minimum $\Delta P=4.5\cdot 10^{-4}$ lies at the true value $c^{\prime}_{6}$. (c) Initial mismatch results in state-dependent purity artifact. We plot test states on the Bloch sphere as circles and color-code their reconstructed purity. (d) When we consider the true retardance, the artifact vanishes.
As another example, we numerically illustrate the measurement-device agnostic calibration for tomography of photonic on-chip path qubits [30, 35, 36, 37]. The qubit is encoded into the photon propagation in one or another waveguide (states $\ket{\uparrow}$ or $\ket{\downarrow}$), as depicted in Figure 3(a). The local heating of one waveguide controls the phase between the two paths. The heating is typically achieved by an electrical current flowing through a small heater element. A Mach-Zehnder interferometer serves as a variable coupler, and its interferometric phase controls the coupling ratio. Detection of photons in a single optical path implements a projective measurement of the path qubit. The circuit in Figure 3(a) implements a projection onto state |$\ket{\pi}$, whose colatitude and longitude on the Bloch sphere are controlled by optical phases. The relation between optical phases $\varphi_{1,2}$ and voltages $V_{1,2}$ applied on the heating elements is approximated with polynomials [35]
@@
Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately. The other imperfections are neglected. We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters. The simulated experiment, depicted in Fig. 3(c), shows the measurement operator mismatch artifact. We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j})$ where the systematic error is mitigated. Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c^{\prime}_{6}$. Taking the true values $c^{\prime}_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
@@ VIII Conclusion
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [13, 22]. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches such as Monte-Carlo sampling [38, 39, 40, 41], permutationally invariant tomography [42], compressed sensing [43, 44], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [45, 46]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [26, 27, 25, 47, 28].
@@ Acknowledgements.
We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grant No. IGA-PrF-2024-008. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.

### v5

@@ IV Special case: multiplicative error
$\tilde{\theta_{j}}=(1+\delta)\theta_{j},$ (6)
$\tilde{\phi_{j}}=(1+\epsilon)\phi_{j}.$ (7)
@@
Because parameters $(\delta,\epsilon)$ are shared among all projection operators, the optimum search simplifies from 12 parameters to just two. In Supplementary material S5, we show that with just eight probe states, the optimization landscape possesses a single minimum, and therefore standard gradient-descent optimization is applicable. We observe that with the less general error model, one can tailor the probe state distribution to achieve the calibration with a smaller probe state ensemble using fewer computational resources. We further illustrate this on an edge example of the phase over-rotation model in Supplementary Material S5.
Another advantage in the case of multiplicative errors is that the global optimum is now unique, and the method unambiguously finds the true set $\{\hat{\pi}_{j}\}$ even when there is an unknown unitary operation between preparation and tomographic measurements. And finally, one can take $\epsilon$ and $\delta$ into account when controlling $\theta_{j}$ and $\phi_{j}$ to achieve a more precise control of the projections.
Having mitigated systematic errors in the measurement devices, we next extend the method to calibrated state preparation, solving the proverbial egg-chicken problem of characterizing quantum measurement without precisely known quantum states.
@@
V State preparation
The state preparation typically suffers from imperfections too. Here, we use the symmetry in the Born rule to formally exchange state preparation and measurement to utilize the previously introduced method to correct state preparation. Usually, the qubit is initialized in state $|0\rangle$, then turned into state $|\psi_{j}\rangle=\hat{U}_{j}|0\rangle$ by unitary evolution $\hat{U}_{j}$. For perfect preparation, $\hat{U}_{j}$ has to be perfectly known. In practice, there might be a slight discrepancy between the desired and the actual evolution. Our calibration reveals this discrepancy and brings information about the actual unitary.
The tomographic projection is also typically realized by some unitary evolution $\hat{V}_{k}$ and subsequent projection, i.e., $|\pi_{k}\rangle=\hat{V}_{k}^{\dagger}|0\rangle$. The state tomogram of the probe state $|\psi_{j}\rangle$ consists of measured probabilities $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle 0|\hat{V}_{k}\hat{U}_{j}|0\rangle|^{2}$. Here, we assume that the operator $\hat{V}_{k}$ is precisely known. If not, the calibration procedure described above might be used to improve the knowledge.
Let us now leverage the Born rule symmetry, i.e. $|\langle\hat{\pi}_{k}|\psi_{j}\rangle|^{2}=|\langle\hat{\psi}_{j}|\pi_{k}\rangle|^{2}$, to reverse our calibration from measurements to preparations. Physical reversal of the process consists of preparing states $\hat{V}_{k}|0\rangle$ and projecting them onto states $\hat{U}_{j}^{\dagger}|0\rangle$, obtaining tomogram elements $p_{kj}\propto|\langle 0|\hat{U}_{j}\hat{V}_{k}|0\rangle|^{2}$. We can analyze the tomograms to improve our knowledge of the measurement operator, i.e., $\hat{U}_{j}^{\dagger}|0\rangle\langle 0|\hat{U}_{j}$, related to state preparation in the non-reversed situation.
The reversal can be realized in the experiment without its physical rearrangement, by performing projections onto the original probe states in the tomographic device and preparing states corresponding to the original tomographical projectors instead of the original probes. Either way, we gain information about $\hat{U}_{j}$, increasing the accuracy of state preparation.
As an example, first recall the situation from Section IV. To calibrate six measurement operators, we applied them to eight probe states. These measurements were implemented by unitary rotations $V_{k}$, followed by projection onto the state $|0\rangle$. The calibration effectively revealed information about the $V_{k}$, i.e. parameters $\delta$ and $\epsilon$ which can be used to more precisely control the projections.
Next, one would prepare six new probe states corresponding to the original projection $V_{k}|0\rangle$. These states carry information about systematic errors in the preparation. Each of these would be projected onto the original eight probe states $U_{j}|0\rangle$. The resulting tomograms contain readings corresponding to imprecise preparations measured with precise measurements. It could then be analyzed to gain information about the parameters of the preparation device. This information can be used to correct the state preparation, as we will demonstrate experimentally in the following section.
@@ VI Experiment: polarization state tomography
We experimentally demonstrate the method in the waveplate-based polarization state tomography scenario. At the end of the process, we will improve our knowledge of the waveplate retardance and consequently mitigate the reconstruction artifact.
The experimental setup is depicted in Figure 2(a). We used a continuous-wave 810 nm thermally-stabilized fiber-coupled laser diode (LD). The laser beam was decoupled into free space using an 11-mm aspheric lens and horizontally polarized using a Glan-Taylor prism (GT). The waveplates HWP1 and QWP1 effectively control the prepared polarization state. The following operator describes the action of a waveplate:
@@
where $c(\theta)=\pi/4$ if $\theta>\pi/2$ and $c(\theta)=0$ otherwise. This relation holds for wave plates that have perfect retardance. The deviation from the ideal retardance will introduce errors in the preparation. We assumed the perfect retardance in the state preparation stage as our method does not require exact knowledge of probe states.
@@
To demonstrate the tolerance of the method to a unitary operation between the preparation and tomographic measurement, we add a quarter-wave plate (QWPX) with its fast axis oriented at -12.5 deg relative to the horizontal direction. Because this unitary operation can be in principle unknown, we restrict ourselves from using any information about QWPX in the calibration.
@@
The states were analyzed using half-waveplate and quarter-wave (HWP2, QWP2) and subsequent projection onto horizontal and vertical polarization using a calcite beam-displacer (BD). The waveplate settings for the Pauli tomography are listed in Tab. S3 of the Supplementary Material. All waveplates were mounted into a precision motorized rotation stage that provided bidirectional repeatability of $\pm 30$ mdeg.
@@
We used reverse-biased PIN photodiodes (PD) to measure optical intensity at both outputs of the beam-displacer. We normalized the readings using the sum of the two signals to eliminate the influence of laser power fluctuation.
@@
The action of the waveplate could be seen as the rotation of the Bloch sphere around axis $(\sin(2\alpha),0,\cos(2\alpha))$ with rotation angle $\Gamma$, where $\alpha$ determines the angular position of the waveplate and $\Gamma$ its retardance. The retardance usually deviates from its nominal value. These deviations manifest as under- or over-rotation of the Bloch sphere. So, in the context of the previous sections, the deviation of the effective wave plate retardance from its nominal value is the origin of the mismatch between the measurement operators. The retardance mismatch can be attributed to the manufacturing tolerances or to the use of a light at a slightly different wavelength than the nominal design value of the wave plate. The exact values of these deviations are initially unknown to us, and we will use the presented method to estimate their values more precisely. We parameterize the measurement operator by precisely controlled waveplate angular positions $x$, $y$, and half- and quarter-wave plate retardance deviations $\delta^{\prime}$ and $\epsilon^{\prime}$, respectively.
@@
We sequentially prepare eight distinct polarization states and perform the Pauli tomography for each. As we introduced in previous sections, these probe states have similar purity and are quasi-uniformly distributed over the Bloch sphere. We use them to probe the measurement device. Initially, we assume zero deviations, $\delta^{\prime}=\epsilon^{\prime}=0$, and observe the purity modulation $\Delta P=0.0572(3)$. The number in parentheses is one standard deviation at the last significant digit, e.g. $0.0572(3)=(572\pm 3)\times 10^{-4}$. The uncertainties were estimated using bootstrapping with the assumption of normally distributed photo-current with 0.1% relative standard deviation, which is the most dominant term and is related to the ammeter precision. The shot noise is negligible at the power level of our laser, compared to the technical noise. Then, we vary parameters $\delta^{\prime}$ and $\epsilon^{\prime}$ and update the measurement operators employed in the reconstruction to minimize the purity modulation. The optimal $\delta$ and $\epsilon$ are close to the true values because the reconstruction artifact is significantly reduced to $\Delta P=0.0140(2)$. We found the optimal parameters to be $\delta^{\prime}=4.5(2)$ deg, $\epsilon^{\prime}=-1.30(4)$ deg.
@@
We then reverse the process to determine retardance deviations $\tilde{\delta},\tilde{\epsilon}$ of the waveplates used in the preparation stage. We prepare the six eigenstates of Pauli operators, project each onto eight states corresponding to the original probe state set, and treat this data as before. We minimize the apparent purity modulation from 0.0700(3) to 0.0144(2) to find the optimal parameters $\tilde{\delta}=4.53(6)$ deg and $\tilde{\epsilon}=-3.58(3)$ deg.
@@
The calibration step revealed the retardance deviation of the wave plates used for preparation and projection. To reconstruct the quantum states, we could directly use this knowledge to update our measurement operators. However, we take one extra step, which allows us to prepare the states with greater precision and control the measurement operators precisely. We took these deviations into account in the verification step and calculated new angular positions of wave plates for tomographic projections and state preparation. Since Eq. 10 does not hold anymore, we used numerical optimization to find the desired wave plate settings ($x,y$):
@@
As a reference, we first assumed perfect waveplates and performed state tomography for each test state. We observed purity modulation $\Delta P=0.0607(2)$ with purity minimum $P_{\min}=0.9383(2)$. This modulation indicates reconstruction artifacts. The minimum fidelity of the reconstructed state to the target states was $F_{\min}=0.96688(4)$. These fidelity values indicate imprecise preparation.
In contrast, when we used our knowledge of $\delta$ and $\epsilon$ in the reconstruction and preparation, the minimum purity increased to $0.9830(3)$ and the minimum fidelity to $F_{\min}=0.9907(2)$. The information gained in the calibration clearly improved the reconstruction. The overall fidelity $F=0.997(3)$ surpasses the value 0.99(1) reported previously by [15]. Note that this fidelity is calculated relatively to the expected ideal state and was evaluated on a set of 58 test states, independently of calibration. This is in contrast to the previous demonstration, where only 14 states were used, and these states were also used for calibration.
@@
Having demonstrated experimental feasibility with polarization-encoded optical qubits, we now turn to on-chip photonic qubits and illustrate how our calibration method applies to path-encoded qubits in this increasingly relevant setting.
@@ VII Path-encoded qubits on a photonic chip
Figure 3: Application of the method for an on-chip photonic path qubit. (a) Projection apparatus for a path qubit. Applied voltage $V_{1,2}$ on heater elements (orange) causes thermal-based phase shift in the waveguide interferometers. These shifts determine the measurement projector. The photons in the upper output path are detected with a single-photon detector (SPAD). (b) 1D-section of 6D function $\Delta P(c^{\prime}_{1},\dots,c^{\prime}_{6})$. The value of $\Delta P$ at the initial assumption $c^{\prime}_{j}$ is $1.8\times 10^{-2}$ on the ensemble of probe states. Minimum $\Delta P=4.5\times 10^{-4}$ lies at the true value $c_{6}$. (c) Initial mismatch results in a state-dependent purity artifact. We plot test states on the Bloch sphere as circles and color-code their reconstructed purity. (d) When we consider the true retardance, the artifact vanishes.
[…]
@@
Here, the voltages $V_{1,2}$ are controlled perfectly, but parameters $c_{j}$ are known only approximately. The other imperfections are neglected. We have increased the number of probe states to 22 to compensate for the higher number of optimization parameters. The simulated experiment, depicted in Fig. 3(c), shows the measurement operator mismatch artifact. We search the six-dimensional parametric space to find the point of minimal $\Delta P(c_{j}^{\prime})$ where the systematic error is mitigated. Panel (b) shows a one-dimensional section illustrating that the minimal $\Delta P$ corresponds to the true value $c_{6}$. Taking the true values $c_{j}$ into account mitigates the reconstruction artifact, as shown in panel (d).
@@ VIII Conclusion
The application of the reported measurement-device agnostic quantum measurement goes beyond the full quantum tomography [11, 22]. The same method will find its use for mitigation of measurement errors in various approximative and more scalable approaches, such as Monte-Carlo sampling [39, 40, 41, 42], permutationally invariant tomography [43], compressed sensing [44, 45], etc. The method can also improve the accuracy of building and characterization of the photonic and quantum circuits in complex media [46, 47]. Finally, the reported approach also applies to experimental realizations of entanglement certification and quantification [26, 27, 25, 48, 28].
@@ Acknowledgements.
We acknowledge the support of the Czech Science Foundation under grant No. 21-18545S. M. B. acknowledges the support of Palacký University under grants No. IGA-PrF-2024-008 and IGA-PrF-2025-010. We acknowledge the use of cluster computing resources provided by the Department of Optics, Palacký University Olomouc. We thank J. Provazník for maintaining the cluster and providing support.